OpenStudy (anonymous):

In an attempt to be environmentally friendly, the city council decides that there must be at least 18 trees on each block in the city. A sample of 20 blocks was randomly surveyed, with these results (number of trees per block): 24, 17, 7, 20, 28, 27, 10, 21, 30, 8, 30, 27, 22, 27, 27, 13, 27, 20, 24, 17. What is the sample mean number of trees per block (x¯)? 17.7 21.3 24 there is insufficient data to answer this question

OpenStudy (anonymous):

@imqwerty can you help again?

OpenStudy (anonymous):

@robtobey please help!!

OpenStudy (anonymous):

you add all the numbers and divide it by 20

OpenStudy (anonymous):

426/20 = 213/10 = 21.3

OpenStudy (anonymous):

I got 21.3

OpenStudy (anonymous):

Congratulations. 21.3 must be the correct answer.

OpenStudy (anonymous):

thank you :)